OpenStudy (anonymous):

use the discriminant to determine the nature of the roots of 2*^2=0

OpenStudy (unklerhaukus):

\[ax^2+bx+c=0\]\[\Delta=b^2-4ac\]

OpenStudy (unklerhaukus):

my guess is that you ment to write \(2x^2=0\) in which case \(a=2;\qquad b=0;\qquad c=0\) so the discriminant would be Δ=0

OpenStudy (unklerhaukus):

which is the case of a one root, (or two equal roots)